# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

### v1

@@ 3.2 Model Architecture
As shown in Figure 1, we adopt AM (Kool et al., 2019) as our basic model. While UDR provides cross-problem consistency, different problems have diverse geometric priors. To efficiently learn the geometric and relational biases inherent in various problems and obtain better generalization performance, we propose two key enhancements: (1) Mixed Bias Modules (MBM) that integrate diverse geometric priors, and (2) adaptive parameter generators conditioned on the UDR. Their implementations are detailed below.
@@
Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project each component into $d$-dimensional embeddings through three separate linear transformations:
@@
where $W_{\bm{\rho}}\in\mathbb{R}^{3\times d},\ W_{\bm{\omega}}\in\mathbb{R}^{6\times d},\ W_{\bm{\xi}}\in\mathbb{R}^{5\times d}$ are learnable matrices. Then we obtain a set of initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$ for all nodes in instance $\mathcal{G}$.
Encoder
The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. The detailed process is provided in Appendix E.1. Unlike prior work, we replace the plain attention (Vaswani et al., 2017) with the proposed mixed bias module (MBM) that incorporates three structural bias matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D^{\mathrm{T}}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup–delivery pairing in PDPs) and 1 otherwise, so that smaller values remain preferable. For the $\ell$-th layer, the MBM output $\hat{\mathbf{h}}_{i}^{(\ell)}$ can be expressed as:
@@
$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases},$ (4)
$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\hat{\mathbf{h}}_{i}^{(0)},\hat{\mathbf{h}}_{i}^{(1)},\hat{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrices. For $\mathrm{Attention}(\cdot)$, we adopt the Adaptation Attention Free Module (AAFM) (Zhou et al., 2024a), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details). We generate $\alpha$ via a lightweight bias network BIAS($\bm{\lambda}$) whose input is the instance’s multi-hot representation $\bm{\lambda}$:
@@
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d},\ W_{2}\in\mathbb{R}^{d\times 1},\ \mathbf{b}_{1}\in\mathbb{R}^{d},\ \mathbf{b}_{2}\in\mathbb{R}^{1}$ are learnable parameters. We set the minimum value of the generated bias to $1$, resulting in faster model convergence. Due to the relation $r_{ij}$ is scale-independent, we remove scale $N$ in adaptation function for $\bm{R}$, and if $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$. As shown in Figure 1(b), unlike MatNet (Kwon et al., 2021), which processes $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ using two parallel attention layers, we mix them into one shared attention layer and let separate heads address different priors. This head‑level incorporation not only cuts architectural redundancy but also enhances model flexibility (e.g., integrating relation matrix), enabling URS can achieve impressive performance for various VRP variants.
Decoder
@@
In URS, all parameters of $\bm{\theta}_{dec}(\bm{\lambda})$ are generated based on the multi-hot problem representation $\bm{\lambda}$. Following Kwon et al. (2020), given the first and last visited node embeddings $\mathbf{h}_{\pi_{1}}^{(L)}$ and $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and an optional constraint state $C_{t}\in\mathbb{R}^{1}$ (e.g., remaining load), the context embedding $\mathbf{h}_{(C)}^{t}$ can be expressed as:
where $\cdot$$\cdot$] is the horizontal concatenation operator, $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ is a linear projection matrix conditioned on $\bm{\lambda}$. Missing $C_{t}$ is zero-padded (see Appendix E.3 for detailed settings). The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2024a) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},H^{(L)}W_{K}(\bm{\lambda}),H^{(L)}W_{V}(\bm{\lambda}),\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (8)
where $W_{K}(\bm{\lambda})\in\mathbb{R}^{d\times d}$, $W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$. Finally, we compute the selection probability for each feasible node $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})$ based on the improved compatibility with adaptation bias (Zhou et al., 2024a):
$p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))+\mathcal{M}_{t})$ (9)
where $\zeta$ is the clipping parameter. Inspired by Lin et al. (2022), we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ (see Appendix E.2 for implementation details). For $\alpha$ in Equation 8 and Equation 9, we generate them via $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6. As shown in Figure 2, compared with adapter-based frameworks (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a), we provide a unified decoding entrance, which $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$ maps a problem representation vector $\bm{\lambda}$ directly to all parameters of the decoder. This conditional weight generation removes problem‑specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
4 LLM-driven Constraint Satisfaction Mechanism
Figure 3: The overview of LLM-driven Constraint Satisfaction Mechanism. (a) The generating pipeline of LLM-driven Constraint Satisfaction Mechanism for a general routing problem; (b) The illustration of mask generator utilization, handling the complex constraint.
While the proposed UDR enables URS to handle a wide range of routing problems, ensuring solution feasibility across diverse and potentially unseen constraints remains a major challenge. Traditional NCO approaches rely on manually designed masks $\mathcal{M}_{t}$ for each specific VRP variant Kool et al. (2019); Kwon et al. (2020); Luo et al. (2023), which becomes impractical when solving over 100 distinct variants as targeted by URS. To address this bottleneck, we propose an LLM-driven constraint satisfaction mechanism that automatically produces problem-specific mask generators. This mechanism transforms constraint handling from a manual engineering task into an automated program design task, thereby effectively supporting URS’s excellent zero-shot generalization capabilities.
As illustrated in Figure 3 (a), our constraint satisfaction mechanism operates through a systematic generate-and-validate pipeline. The process begins by prompting an LLM with both a problem-specific natural language constraint description and a standardized Python code template. The LLM then designs a set of candidate mask generators that implement the constraint logic. Each candidate is validated on a set of problem instances using a problem-specific checker to verify its reliability in producing feasible solutions. Invalid candidates are discarded, and the generation process iterates until a reliable mask generator is obtained.
Specifically, during validation, the candidate mask generator guides the stepwise solution construction. As illustrated in Figure 3 (b), at each step $t$, the mask generator takes the current state of the partial solution $\pi_{1:t-1}$ as input, which contains various information such as visited nodes, remaining load, etc. It then produces a binary mask $\mathcal{M}_{t}$ that identifies all permissible nodes. This mask is directly integrated into the URS’s selection probability computation in Equation 9 to ensure feasibility.
To efficiently steer this generate-and-test process, we formulate it as an optimization problem that maximizes the solution validity rate. The well-developed LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), can be effectively applied to tackle this optimization problem. The optimization process terminates immediately once a reliable mask generator is found, which significantly reduces computational overhead compared to exhaustive search approaches. Furthermore, generated mask generators can be cached and reused for different instances sharing the same problem variant, thereby amortizing the generation cost. For more details about our LLM-driven constraint satisfaction mechanism, please see Appendix G.
@@ 5 Experiments
In this section, we comprehensively evaluate URS against classical and learning-based solvers across more than one hundred distinct VRP variants. For URS, we focus on: (1) performance on seen problems and (2) zero-shot generalization under unseen ones. To ensure comparability with prior neural routing studies, we adopt the 100-node setting as our primary evaluation scenario for URS. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory).

### v2

@@ 3.2 Model Architecture
As shown in Figure 1, we adopt AM (Kool et al., 2019) as our basic model. While UDR provides cross-problem consistency, different problems have diverse geometric priors. To efficiently learn the multiple priors inherent in various problems and obtain better cross-problem zero-shot generalization, we propose two key enhancements: (1) a Mixed Bias Module (MBM) that captures multiple priors inherent to various problems, and (2) problem-conditioned parameter generators conditioned on the UDR. Their implementations are detailed below.
@@
Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project it into $d$-dimensional embeddings through three separate linear transformations:
@@
where $W_{\bm{\rho}}\in\mathbb{R}^{3\times d},\ W_{\bm{\omega}}\in\mathbb{R}^{6\times d},\ W_{\bm{\xi}}\in\mathbb{R}^{5\times d}$ are learnable matrices. Then we obtain a set of initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$ for all nodes in instance $\mathcal{G}$. This embedding is the initial input of encoder $\bm{\theta}_{enc}$. After passing through $L$ stacked attention layers, $H^{(0)}$ is transformed into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. The detailed encoding process is provided in Appendix E.1.
Figure 2: The structure of the proposed Mixed Bias Module, explained on a 44-node APDTSP instance. We replace plain attention with MBM during encoding to improve node embeddings, which effectively captures multiple priors inherent to various problems while reducing architectural redundancy.
Mixed Bias Module
Given the distinct priors across VRP variants, efficiently capturing the intrinsic geometric bias is pivotal for enhancing the cross-problem generalization of RL-based multi-task neural solvers. Current leading RL-based models generally adopt a heavy encoder and light decoder architecture, where the quality of the encoder-generated node embeddings plays a key role in overall performance. While prior studies like MatNet (Kwon et al., 2021) have attempted to handle asymmetry by integrating directional distance biases via parallel layers, such designs incur high model complexity and lack universality. Consequently, effectively encoding multiple inherent biases, such as symmetric and asymmetric distances, and relational dependencies, remains challenging, hindering the acquisition of high-quality solutions across a wide range of problems.
To address the limitation, we propose the mixed bias module (MBM) to replace the vanilla attention mechanism (Vaswani et al., 2017), which enhances cross-problem generalization while reducing architectural redundancy. As shown in Figure 2, MBM can efficiently capture three problem-specific prior matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D}^{\mathrm{T}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup-delivery pairing) and 1 otherwise. This ensures that smaller values are preferred, aligning with the distance metric, where smaller values correspond to a larger selection bias. For the $\ell$-th layer, the MBM output $\hat{\mathbf{h}}_{i}^{(\ell)}$ can be expressed as:
@@
$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases}$ (4)
$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\bar{\mathbf{h}}_{i}^{(0)},\bar{\mathbf{h}}_{i}^{(1)},\bar{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrix. For $\mathrm{Attention}(\cdot)$, we adopt the attention mechanism introduced by Zhou et al. (2026), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details). Note that since the relation $r_{ij}$ is scale-independent, we remove the scale $N$ in the adaptation function for $\bm{R}$. If $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$. For a comparison between our MBM and existing attentions (Vaswani et al., 2017; Kwon et al., 2021), please see Appendix G.
Problem-Conditioned Parameter Generation
Using a single static set of parameters to solve 110 VRP variants is a significant challenge. To mitigate this, we introduce a problem-conditioned mechanism based on problem representation $\bm{\lambda}$, which consists of two components: (1) $\mathrm{BIAS}(\bm{\lambda})$ adjusts the degree to which different priors are integrated into the attention; and (2) $\mathrm{WEIGHT}(\bm{\lambda})$ generates decoder parameters for each problem.
As shown in Figure 3, unlike existing adapter-based fine-tuning methods (Lin et al., 2024; Drakulic et al., 2025), we provide a unified decoding entrance and enable URS to adaptively generate decoder parameters for each problem. This conditional weight generation removes problem-specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
For MBM in Equation 5, we generate $\alpha$ via a lightweight bias network $\mathrm{BIAS}(\bm{\lambda})$:
@@
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d},\ W_{2}\in\mathbb{R}^{d\times 1},\ \mathbf{b}_{1}\in\mathbb{R}^{d},\ \mathbf{b}_{2}\in\mathbb{R}^{1}$ are learnable parameters. We set the minimum bias to $1$, resulting in faster model convergence.
For decoder $\bm{\theta}_{dec}(\bm{\lambda})$, all parameters are generated based on $\bm{\lambda}$. Following Kwon et al. (2020), given the first and last visited node embeddings $\mathbf{h}_{\pi_{1}}^{(L)}$ and $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and an optional constraint state $C_{t}\in\mathbb{R}^{1}$ (e.g., remaining load), the context embedding $\mathbf{h}_{(C)}^{t}$ can be expressed as:
where $\cdot$$\cdot$] is the horizontal concatenation operator, $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ is a linear projection matrix conditioned on $\bm{\lambda}$. Missing $C_{t}$ is zero-padded (see Appendix E.3 for details). The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2026) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
$K=H^{(L)}W_{K}(\bm{\lambda}),\qquad V=H^{(L)}W_{V}(\bm{\lambda}),$ (8)
$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},K,V,\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (9)
where $W_{K}(\bm{\lambda}),W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$. Finally, we compute probabilities for feasible nodes $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\mathbf{u})$ based on the improved compatibility (Zhou et al., 2026):
$u^{t}_{i}=\begin{cases}\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))&\text{if}\ i\not\in\{\pi_{1:t-1}\}\\ -\infty&\text{otherwise,}\end{cases}$ (10)
@@
where $\zeta$ is the clipping parameter. Inspired by Lin et al. (2022), we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ (see Appendix E.2 for implementation details). For $\alpha$ in Equation 9 and Equation 10, we generate them via $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6.
Table 2: Performance comparison on $8$ seen less-explored variants.
Method OP100 PCTSP100 PDTSP100 ACVRP100 CVRPTW100 CVRPB100 OCVRP100 OCVRPTW100
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m 0.00% 6.6m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
Sym-NCO 0.68% 15s 0.14% 16s - - - - - -
BQ 0.63% 7s - - - - - - -
Heter-AM - - 6.61% 5m - - - - -
MVMoE - - - - 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
MTPOMO - - - - 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
ReLD-MTL - - - - 4.56% 10s 0.90% 6s 2.32% 7s 3.10% 9s
GOAL-MTL 1.20% 38s - - - 4.66% 42s - - -
URS-STL 0.11% 4s -0.09% 5s 3.22% 4s 2.01% 1.4m 4.50% 8s 1.59% 6s 2.25% 6s 3.22% 8s
URS 0.45% 4s 1.06% 5s 4.98% 4s 3.06% 1.4m 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
Table 3: Experimental results on seen ATSP, TSP, and CVRP.
Method ATSP100 TSP100 CVRP100
Gap Time Gap Time Gap Time
Oracle 0.00% 2m 0.00% 6m 0.00% 9.1m
Sym-NCO - 0.14% 5s 1.46% 6s
BQ 1.27% 1m 0.35% 7s 3.24% 8s
LEHD - 0.59% 2s 3.68% 2s
POMO - 0.13% 5s 1.25% 6s
ICAM 3.05% 1m 0.15% 4s 2.04% 4s
MatNet 0.95% 6.5m - -
MVMoE - - 1.65% 12s
MTPOMO - - 1.85% 6s
ReLD-MTL - - 1.42% 8s
GOAL-MTL 1.77% 1m - 4.22% 48s
URS-STL 0.62% 1.1m 0.08% 6s 1.63% 6s
URS 2.26% 1.1m 0.57% 6s 1.81% 6s
@@ 4 Experiments
In this section, we present a comprehensive evaluation of URS against both classical and neural solvers. For URS, we focus on four key aspects: (1) performance on 11 seen problems; (2) zero-shot generalization on 99 unseen problems; (3) scalability to large-scale VRP variants; and (4) results on benchmark datasets. All experiments are conducted using a single NVIDIA GeForce RTX 4090 GPU (24GB of memory) for neural solvers and an Intel Xeon Gold 6240R CPU (2.40 GHz) for classical solvers.
Table 4: Zero-shot generalization on 1K test instances of $32$ unseen VRPs ($N=100$). Due to page limit, complete experimental results (including all 99 unseen variants) are deferred to Appendix I.
Method CVRPBP CVRPBPL MDOCVRPBP MDOCVRPBPTW MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 13.38% 14s 56.60% 33s 63.77% 38s 56.26% 34s 62.47% 40s 33.06% 31s 29.08% 31s
MTPOMO 13.71% 7s 13.44% 8s 42.94% 24s 47.09% 29s 41.98% 27s 46.26% 32s 23.38% 23s 27.46% 23s
ReLD-MTL 13.57% 9s 12.96% 10s 37.74% 27s 45.16% 34s 37.45% 29s 44.74% 36s 18.39% 25s 22.32% 25s
URS 12.95% 7s 12.65% 8s 24.44% 22s 26.31% 26s 24.22% 24s 26.20% 28s 15.05% 23s 22.01% 23s
Method MDCVRPL MDOCVRPTW MDOCVRPB MDOCVRPBL MDOCVRPLTW SPCTSP PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h 0.00% 1.6h 0.00% 1.6h
MVMoE 33.42% 34s 51.57% 40s 28.21% 26s 28.70% 28s 51.09% 43s - - -
MTPOMO 24.32% 25s 35.71% 29s 25.06% 20s 25.01% 21s 35.39% 31s - - -
ReLD-MTL 18.48% 28s 33.55% 33s 18.50% 21s 18.50% 22s 33.26% 35s - - -
URS 15.32% 26s 26.05% 25s 15.17% 18s 14.90% 20s 25.59% 27s -2.37% 5s -1.47% 5s 4.93% 5s
Method ACVRPL ACVRPBL ACVRPB ACVRPBTW APDTSP AMDCVRPL AMDCVRP APDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS -2.24% 1.9m 6.18% 2.4m 7.10% 2.1m 9.95% 2.2m 6.21% 1m 7.15% 6.1m 8.11% 5.3m 7.03% 1.2m
Method ACVRPLTW ACVRPBLTW ACVRPBPLTW AOCVRPBLTW ACVRPTW ACVRPBPTW AOPDCVRP AOP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 12.22% 2.7m 10.18% 2.4m 13.82% 2.8m 14.32% 2.5m 11.28% 2.5m 14.52% 2.6m 11.13% 1.2m -5.47% 1.2m
Table 5: Generalization on different large-scale VRP variants. All models are trained on instances of size $N=100$ and employ instance augmentation $\times$ 8. PyVRP and OR-Tools can process all instances of a given size concurrently to ensure a fair comparison.
Problem Method N = 1000 N = 2000 N = 3000 N = 4000 N = 5000
Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time
CVRPTW PyVRP 159.35(0.00%) 20m 337.15(0.00%) 40m 516.97(0.00%) 1h 618.58(0.00%) 2.7h 787.25(0.00%) 3.3h
OR-Tools 178.13(11.79%) 20m 357.18(5.94%) 40m 533.97(3.29%) 1h 647.92(4.74%) 2.7h 807.128(2.52%) 3.3h
MTPOMO 219.22(37.57%) 1.7m 488.40(44.86%) 14.6m 773.00(49.53%) 49.5m 975.61(57.72%) 1.9h OOM
MVMoE 233.53(46.55%) 1.7m 624.73(85.30%) 14.6m 1094.25(111.67%) 51.9m 1486.34(140.28%) 2h OOM
ReLD-MTL 183.31(15.04%) 1.6m 396.28(17.54%) 13.4m 613.58(18.69%) 45.7m 748.65(21.03%) 1.8h OOM
URS 172.58(8.30%) 1.3m 365.27(8.34%) 11.1m 559.03(8.14%) 30.1m 673.63(8.90%) 1.2h 853.79(8.45%) 2.3h
CVRPB PyVRP 36.52(0.00%) 20m 51.80(0.00%) 40m 75.12(0.00%) 1h 93.52(0.00%) 2.7h 112.21(0.00%) 3.3h
OR-Tools 44.23(21.11%) 20m 63.33(22.26%) 40m 88.6(17.94%) 1h 108.82(16.36%) 2.7h 128.11(14.17%) 3.3h
MTPOMO 45.92(25.77%) 1.1m 83.50(61.20%) 8.8m 136.52(81.73%) 31.4m 197.99(111.70%) 1.2h OOM
MVMoE 98.60(170.03%) 1m 250.47(383.57%) 8.1m 332.93(343.19%) 28.9m 423.19(352.49%) 1.1h OOM
ReLD-MTL 38.52(5.48%) 1m 59.21(14.30%) 8m 88.36(17.62%) 28.3m 114.88(22.83%) 1.1h OOM
URS 34.86(-4.54%) 40s 50.39(-2.72%) 5.1m 72.01(-4.15%) 17.3m 91.74(-1.91%) 41.2m 111.93(-0.24%) 1.3h
OCVRPTW PyVRP 90.91(0.00%) 20m 164.00(0.00%) 40m 224.16(0.00%) 1h 299.45(0.00%) 2.7h 367.86(0.00%) 3.3h
OR-Tools 105.98(16.58%) 20m 207.041(26.24%) 40m 280.88(25.30%) 1h 370.85(23.84%) 2.7h 447.02(21.52%) 3.3h
MTPOMO 147.14(61.85%) 1.6m 297.53(81.43%) 12.8m 437.73(95.28%) 44.3m 601.63(100.91%) 1.7h OOM
MVMoE 147.35(62.09%) 1.6m 389.60(137.57%) 12.9m 634.11(182.88%) 44.7m 899.18(200.28%) 1.7h OOM
ReLD-MTL 110.00(21.01%) 1.5m 208.60(27.20%) 12.3m 292.70(30.57%) 42m 394.42(31.72%) 1.6h OOM
URS 98.76(8.63%) 1m 178.07(8.58%) 7.6m 243.88(8.80%) 25.4m 325.64(8.74%) 1h 398.72(8.39%) 1.9h